# Physics Informed Generative Machine Learning for Accelerated Quantum-centric Supercomputing

### v2

@@ V Results and Discussion
In this section, we numerically demonstrate the accuracy and efficiency of PIGen-SQD against SQD. To assess the performance of our approach we compute the ground state energies of three strongly correlated systems, (equilibrium O-H bond distance , bond angle ), () and () for different geometries using 6-31G basis set. With the frozen core approximation, , and consists of 24, 32 and 40 spin orbitals respectively which are mapped directly to 24, 32 and 40 qubits on IBM Heron R2 quantum processors as per the details discussed in Section IV. For Gibbs sampling process, the number of generated determinants is fixed throughout the iteration process at of the symmetry space dimension . The value is used in all the studies, unless explicitly mentioned otherwise.
@@ V.1 RBM Driven Sampling v/s Random Sampling
To numerically illustrate the degree of anchoring effect of the physics-informed initialization provided to hardware samples, Fig. 4 shows a comparative study regarding different rank of perturbative measures having different initialization to the fermionic state reconstruction process for different geometries of and . This analysis incorporates both symmetry-preserved hardware samples and perturbatively generated initializations with determinants up to ranks . Approximate perturbative estimates at each rank were obtained by including a subset of symbolically generated algebraic terms (see Appendix A, B). The plot demonstrates that ML-based configuration recovery converges much faster with better energy accuracy when initialized with excited configurations up to rank 4 compared to only hardware, (up to) rank 2 and rank 3 initializations. This is observed for both equilibrium and stretched geometries where strong correlation effects are more prevalent. The inset bar plots show the number of configurations required in the search space to reach chemical accuracy. In equilibrium geometries, determinants taken up to rank 4 provide an excellent starting point and hence requires substantially reduced dimensional search space to reach chemical accuracy. This ensures an effective learning by the generative model of the distribution of the target state basis vectors. For stretched geometries also, similar trend persists, though the relative difference between the search spaces narrows down compared to equilibrium geometries (see the inset bar plots) as the presence of strong correlation requires a substantial number of even higher rank determinants and more search space. However, even in very strongly correlated regime where the low order perturbative estimates may miss out a few important configurations, the generative learning algorithm is eventually able to pick those up, albeit with a slower macro-cyclic convergence and somewhat larger search space. Moreover, initializations with raw symmetry-preserved hardware samples show similar accuracy and search space dimensions compared to rank 2 and rank 3 initializations in most of the cases under consideration in Fig. 4. This suggests that one needs to strike the right balance between classical physics-informed initialization and the generative procedure to an accurate and efficacious energy estimation. Toward this, from here onward, we will use for all PIGen-SQD calculations. However, one may choose a higher value of (with some selected algebraic terms that scale favorably) for more strongly correlated systems to accelerate convergence which would be explored further in future works.
@@ V.3 Potential Energy Profile and State-Fidelity:
This section demonstrates the accuracy and efficiency of PIGen-SQD against conventional SQD by comparing the performance of both the methods to predict ground state energies of , and at different bond distances.
@@
The dissociation curve of via symmetric bond stretching in 6-31G basis set with frozen core approximation (8 electrons in 12 spatial orbitals) is shown in Fig. 5 for PIGen-SQD, SQD and other reference methods such as HF, MP2, CISD and CCSD for comparison. Panel (A) of Fig.5 shows the potential energy surface and the associated energy difference in with FCI is shown in panel (B) along y-axis with the ratio of bond distance to equilibrium distance plotted along x-axes. The comparative study of the diagonalization dimension for SQD and PIGen-SQD is shown in panel (C) along with the associated search space, spin-particle number conserved symmetry space and core space dimension. SQD is performed in this case with sample size 500 and 1000 in 5 batches at each step of configuration recovery loop while PIGen-SQD results are obtained with ( here) of the symmetry space samples generated in each macro-cycle. For SQD, the minimum energy and maximum diagonalization dimension among all batches are plotted. SQD with sample size 500 or SQD(500) shows chemically accurate results near equilibrium geometry while it struggles for stretched bonds. Since increasing sample size includes more configurations in the diagonalization subspace, SQD(1000) performs well for the entire dissociation curve. On the other hand, PIGen-SQD() and PIGen-SQD() both shows results well within the chemical accuracy. However, the main aspect that distinguishes the efficiency of SQD and PIGen-SQD is the associated diagonalization dimension. In this study, while SQD(1000) requires almost around 84% of the entire symmetry space for diagonalization to provide chemically accurate results for the entire potential energy surface, PIGen-SQD(2%) does so using only of the same. Moreover, our method is systematically improvable by increasing during generative process as demonstrated in the plot with the PIGen-SQD(6%) that shows substantially improved accuracy compared to SQD(1000) throughout the dissociation curve particularly in the strongly correlated regime. Another important metric of our method is the search space that RBM explores during the self-consistent configuration recovery process. From Fig.5 it is evident that the search space for PIGen-SQD is around the core space dimension for for the entire dissociation curve while it slightly increases for providing better accuracy. However, in all the cases, even the search space dimension is well below that of SQD diagonalization dimension, showing an effective prediction of the underlying distribution of the dominant states in the Hilbert space. While PIGen-SQD can be systematically improved by tuning , in our subsequent studies, we will use as it strikes an optimal balance between the accuracy and the explored Hilbert space.
@@
Bond stretching of the () molecule is a paradigmatic approach to demonstrate electronic strong correlation effects. Fig. 6 shows the dissociation curve of for both SQD and PIGen-SQD along with other standard methods in 6-31G basis with frozen core approximation resulting in 10 electrons in 16 spatial orbitals. The SQD results are obtained using sample size 3000 and 5000 with fixed batch size 5. For some stretched geometries CCSD did not converge and those points are not shown in the figure. While both SQD(3000) and SQD(5000) shows chemically accurate results near equilibrium geometries, it fails in stretched geometries having errors of the order of and with around 6 and 10 million determinants in the diagonalization space respectively. Contrarily, PIGen-SQD(2%) shows drastically improved results maintaining around accuracy throughout the dissociation curve, even under strong correlation at stretched geometries. Remarkably, PIGen-SQD needs only about 1-1.5 million determinants in the diagonalization subspace, which shows almost an order-of-magnitude reduction in the diagonalization dimension while providing more than an order-of-magnitude energy accuracy. The associated search space is well under the core space dimension near equilibirum geometries, while it increases to of the order of core space dimension for stretched geometries. However, the search space is still well below the diagonalization dimension of SQD(5000).
@@
Fig. 7 illustrates an approximate estimation of fidelity of the reconstructed wavefunctions by comparing the CI coefficients obtained from the different methods against FCI coefficients. The absolute CI coefficients obtained from FCI, SQD and PIGen-SQD are sorted in an ascending order and plotted along the x-axis for the two geometries of ((A), (B)) and ((C), (D)). For any given interval on the x-axis, the corresponding y-value indicates the number of coefficients falling within that range. The corresponding energy difference from FCI is given in the inset text for all the cases. For both SQD and PIGen-SQD shows similar order of accuracy though the number of configurations predicted by SQD is substantially large. Fig.7 (A, B) histograms show PIGen-SQD distribution has a peak toward the larger CI coefficients () representing a dominant subset of the FCI coefficients that provides major support to target state and contribute more toward accurate energy estimation. Better state fidelity for PIGen-SQD is more conspicuous for stretched geometries of in Fig. 7 (C, D) as one can see the substantially higher peak of PIGen-SQD curve toward the region with larger CI coefficients than SQD which has a flatter distribution over the entire range of values. This is reflected in the corresponding energy errors (see the inset text) where PIGen-SQD shows more than one order of magnitude better energy accuracy (0.3 compared to 4 obtained by SQD) under strong correlation. These examples demonstrate that PIGen-SQD faithfully reconstructs accurate high-fidelity states as compared to SQD leading to significantly improved energy predictions with substantially reduced requirements of HPC resources. The corresponding density plot in Fig.9 (see Appendix C) shows the predicted coefficients are sharply peaked around larger values with maximum probability for PIGen-SQD, making it a unimodal distribution. On the contrary, SQD and FCI curves are bimodal with a higher peak near practically zero coefficients due to sparsity. This further confirms the concentrated sampling of the dominant sector of the Hilbert space enabled by PIGen-SQD.
@@
: Fig. 8 shows PIGen-SQD versus SQD convergence for two isolated single point calculations of (C-C equilibrium bond distance ). SQD configuration recovery calculations are performed in this case using sample size 5000 in 5 batches. Each column in Fig.8 shows the energy difference with HCI in log-scale (top row) and the number of configurations required (bottom row) in diagonalization subspace and search space of the respective methods against a common x-axis representing number of iterations. Column (A) and (B) shows calculations for C-C bond-length near equilibrium geometry and stretched geometry . PIGen-SQD convergence behaviour shows iteratively improved energy with more than one-order-of-magnitude better accuracy compared to SQD while keeping the diagonalization dimension and search space as low as possible. Quantitatively, even with a 75 reduction in the associated diagonalization subspace dimension for both the cases under consideration, energy accuracy of around 0.1 and 1 is achieved for equilibrium and stretched geometries compared to 5 and 10 for SQD, showing the superiority of PIGen-SQD under strong correlations.
@@ VI Conclusion and Future Outlook
In this paper, we propose PIGen-SQD, a novel QCSC workflow based on SQD framework that leverages a strategic orchestration of sampling efficiency of quantum computers, many-body intuitions and generative power of machine learning to provide better convergence with substantially reduced classical resource utilization. QCSC/ SQD utilizes the quantum computer as a sampling engine to generate configuration subspace in which the molecular Hamiltonian is projected and diagonalized. In the presence of quantum hardware noise, SQD requires a configuration recovery module, a heuristic technique that “corrects” the noise-affected samples with iterative filtering based upon the correct spin-particle number symmetry and average occupation probability. Accuracy of SQD relies upon the the degree of initial overlap with the target state provided by the sampled distribution and the capability of heuristic configuration recovery scheme to predict dominant configurations. However, the key challenge is that the state preparation on current quantum hardware does not necessarily guarantee high quality samples to provide this support to the target state and may require prohibitively huge number of shots which collectively adds a heuristic component to the SQD pipeline. To address this, here we introduce a physics-informed pre-processing which employs perturbative measures with implicit low rank tensor decompositions to identify a particular group of dominant configurations. These perturbative measure when combined with hardware samples, yields a substantially improved initial overlap to the target state providing a head-start for configuration recovery. This is followed by integration of generative ML models (RBM in this case) which upon training with this sample distribution, generates new configurations while exploring only the relevant sector of the Hilbert space of chemical relevance. Our numerical demonstrations with 24, 32 and 40 qubit molecular simulations conducted on state-of-the-art IBM Heron R2 processors suggest PIGen-SQD outperforms traditional SQD in both accuracy and efficiency, providing up to one-order of better energy accuracy with up to 90 reduction in the corresponding subspace dimension required for Hamiltonian diagonalization. Such accuracy with substantially reduced HPC resource requirements is observed even under strong electronic correlations where SQD struggles to predict chemically accurate results.
@@
From a near-future application perspective PIGen-SQD can be integrated with other SQD variants such as Sample-based Krylov Quantum Diagonalization (SKQD)[44] or its more efficient variant using quantum stochastic drift protocol (SqDRIFT)[47] for guaranteed and faster convergence along with enhanced computational efficiency. Moreover, the current architecture have huge scope of technical improvements such as RBM can be replaced with more advanced neural network models such as transformers[72] which has recently been shown to provide unique advantages while dealing with complex molecular systems. With continued improvements in quantum-hardware capabilities and machine-learning components, PIGen-SQD presents a promising pathway toward scalable molecular simulations, advancing the field toward resolving several longstanding challenges in chemistry.

### v3

@@ V Results and Discussion
In this section, we numerically demonstrate the accuracy and efficiency of PIGen-SQD against SQD. To assess the performance of our approach we compute the ground state energies of three strongly correlated systems, (equilibrium O-H bond distance , bond angle ), () and () for different geometries using 6-31G basis set. In addition, we consider in the cc-pVDZ basis to assess the performance of PIGen-SQD with SQD with more quantum hardware resources. For Gibbs sampling process, the number of generated determinants is fixed throughout the iteration process at of the symmetry space dimension . The value is used in all the studies, unless explicitly mentioned otherwise.
@@ V.1 RBM Driven Sampling v/s Random Sampling
To numerically illustrate the degree of anchoring effect of the physics-informed initialization provided to hardware samples, Fig. 3 shows a comparative study regarding different rank of perturbative measures having different initialization to the fermionic state reconstruction process for two different geometries of and in 6-31G basis. This analysis incorporates both symmetry-preserved hardware samples and perturbatively generated initializations with determinants up to ranks . Approximate perturbative estimates at each rank were obtained by including a subset of symbolically generated algebraic terms (see Appendix A, B). The plot demonstrates that ML-based configuration recovery converges much faster with better energy accuracy when initialized with excited configurations up to rank 4 compared to only hardware, (up to) rank 2 and rank 3 initializations. This is observed for both equilibrium and stretched geometries where strong correlation effects are more prevalent. The inset bar plots show the number of configurations required in the search space to reach chemical accuracy. In equilibrium geometries, determinants taken up to rank 4 provide an excellent starting point and hence requires substantially reduced dimensional search space to reach chemical accuracy. This ensures an effective learning by the generative model of the distribution of the target state basis vectors. For stretched geometries also, similar trend persists, though the relative difference between the search spaces narrows down compared to equilibrium geometries (see the inset bar plots) as the presence of strong correlation requires a substantial number of even higher rank determinants and more configurations in the search space. However, even in very strongly correlated regime where the low order perturbative estimates may miss out some important configurations, the generative learning algorithm is eventually able to pick those up, albeit with a slower macro-cyclic convergence and somewhat larger search space. Moreover, initializations with raw symmetry-preserved hardware samples show similar accuracy and search space dimensions compared to rank 2 and rank 3 initializations in most of the cases under consideration in Fig. 3. The benefit of augmenting hardware-derived samples with perturbatively generated configurations is illustrated in Fig. 3(D). In this case for the stretched geometry of , initialization with raw hardware samples (symmetry-preserved) followed by RBM-driven sampling fails to reach chemical accuracy, while PIGen-SQD initialized with various ranks of perturbatively selected configurations reaches desired accuracy consistently. This behavior can be attributed to the severely limited shot budget in this case (), compared to the associated Hilbert-space dimension (), which leads to a poor-quality measurement distribution with negligible overlap with the true ground state. As a result, the RBM is unable to reliably identify the underlying structure of the wavefunction, generating only a small number of dominant determinants that are insufficient for accurate energy estimation. This example highlights the critical role of effective many-body inspired initialization in classical post-processing and demonstrates that, even under stringent shot constraints, PIGen-SQD can efficiently recover dominant configurations when appropriately initialized. This suggests that one needs to strike the right balance between classical physics-informed initialization and the generative procedure to an accurate and efficacious energy estimation. Toward this, from here onward, we will use for all PIGen-SQD calculations. However, one may choose a higher value of (with some selected algebraic terms that scale favorably) for more strongly correlated systems to accelerate convergence which would be explored further in future works.
@@ V.3 Potential Energy Profile and State-Fidelity:
This section demonstrates the accuracy and efficiency of PIGen-SQD against conventional SQD by comparing the performance of both the methods to predict ground state energies of and in 6-31G basis across different geometries.
@@
The dissociation curve of via symmetric bond stretching in 6-31G basis set with frozen core approximation (8 electrons in 12 spatial orbitals) is shown in Fig. 4 for PIGen-SQD, SQD and other reference methods such as HF, MP2, CISD and CCSD for comparison. Panel (A) of Fig.4 shows the potential energy surface and the associated energy difference in with FCI is shown in panel (B) along y-axis with the ratio of bond distance to equilibrium distance plotted along x-axes. The comparative study of the diagonalization dimension for SQD and PIGen-SQD is shown in panel (C) along with the associated search space, spin-particle number conserved symmetry space and core space dimension. SQD is performed in this case with sample size 500 and 1000 in 5 batches at each step of configuration recovery loop while PIGen-SQD results are obtained with ( here) of the symmetry space samples generated in each macro-cycle. For SQD, the minimum energy and maximum diagonalization dimension among all batches are plotted. SQD with sample size 500 or SQD(500) shows chemically accurate results near equilibrium geometry while it struggles for stretched bonds. Since increasing sample size includes more configurations in the diagonalization subspace, SQD(1000) performs well for the entire dissociation curve. On the other hand, PIGen-SQD() and PIGen-SQD() both shows results well within the chemical accuracy. However, the main aspect that distinguishes the efficiency of SQD and PIGen-SQD is the associated diagonalization dimension. In this study, while SQD(1000) requires almost around 84% of the entire symmetry space for diagonalization to provide chemically accurate results for the entire potential energy surface, PIGen-SQD(2%) does so using only of the same. Moreover, our method is systematically improvable by increasing as demonstrated in the plot with the PIGen-SQD(6%) that shows substantially improved accuracy compared to SQD(1000) throughout the dissociation curve particularly in the strongly correlated region. Another important metric of our method is the search space that RBM explores during the self-consistent configuration recovery process. From Fig.4 it is evident that the search space for PIGen-SQD is around the core space dimension for for the entire dissociation curve while it slightly increases for providing better accuracy. However, in all the cases, even the search space dimension is well below that of SQD diagonalization dimension, showing an effective prediction of the underlying distribution of the dominant states in the Hilbert space. While PIGen-SQD can be systematically improved by tuning , in our subsequent studies, we will use as it strikes an optimal balance between the accuracy and the explored Hilbert space.
@@
Bond stretching of the () molecule is a paradigmatic approach to demonstrate electronic strong correlation effects. Fig. 5 shows the dissociation curve of for both SQD and PIGen-SQD along with other standard methods in 6-31G basis with frozen core approximation resulting in 10 electrons in 16 spatial orbitals. The SQD results are obtained using sample size 3000 and 5000 with fixed batch size 5. For some stretched geometries CCSD did not converge and those points are not shown in the figure. While both SQD(3000) and SQD(5000) shows chemically accurate results near equilibrium geometries, it fails in stretched geometries having errors of the order of and with around 6 and 10 million determinants in the diagonalization space respectively. Contrarily, PIGen-SQD(2%) shows drastically improved results maintaining around accuracy throughout the dissociation curve, even under strong correlation at stretched geometries. Remarkably, PIGen-SQD needs only about 1-1.5 million determinants in the diagonalization subspace, which shows almost an order-of-magnitude reduction in the diagonalization dimension while providing more than an order-of-magnitude energy accuracy. The associated search space is well under the core space dimension near equilibirum geometries, whereas for stretched geometries it grows to be comparable in size to the core-space dimension. However, the search space is still well below the diagonalization dimension of SQD(5000).
@@
Fig. 6 illustrates an approximate estimation of fidelity of the reconstructed wavefunctions by comparing the CI coefficients obtained from the different methods against FCI coefficients. The absolute CI coefficients obtained from FCI, SQD and PIGen-SQD are sorted in an ascending order and plotted along the x-axis for the two geometries of ((A), (B)) and ((C), (D)). For any given interval on the x-axis, the corresponding y-value indicates the number of coefficients falling within that range. The corresponding energy difference from FCI is given in the inset text for all the cases. For , both SQD and PIGen-SQD shows similar order of accuracy though the number of configurations predicted by SQD is substantially large. Fig.6 (A, B) histograms show PIGen-SQD distribution has a peak toward the larger CI coefficients () representing a dominant subset of the FCI coefficients that provides major support to target state and contribute more toward accurate energy estimation. Better state fidelity for PIGen-SQD is more conspicuous for stretched geometries of in Fig. 6 (C, D) as one can see the substantially higher peak of PIGen-SQD curve toward the region with larger CI coefficients than SQD which has a flatter distribution over the entire range of values. This is reflected in the corresponding energy errors (see the inset text) where PIGen-SQD shows more than one order of magnitude better energy accuracy (0.1 compared to 4 obtained by SQD) under strong correlation. These examples demonstrate that PIGen-SQD faithfully reconstructs accurate high-fidelity states as compared to SQD leading to significantly improved energy predictions with substantially reduced requirements of HPC resources. The corresponding density plot in Fig.9 (see Appendix C) shows the predicted coefficients are sharply peaked around larger values with maximum probability for PIGen-SQD, making it a unimodal distribution. On the contrary, SQD and FCI curves are bimodal with a higher peak near practically zero coefficients due to sparsity. This further confirms the concentrated sampling of the dominant sector of the Hilbert space enabled by PIGen-SQD.
V.4 Beyond Exact Diagonalization
Since the Hilbert space grows exponentially with system size, exact diagonalization becomes infeasible for larger molecular systems and extended basis sets. We therefore consider two representative cases in the 6-31G basis and in the cc-pVDZ basis for which exact diagonalization is computationally impractical. In these regimes, we assess the performance of PIGen-SQD in comparison with standard SQD.
@@ V.4 Beyond Exact Diagonalization
Fig. 7 shows PIGen-SQD versus SQD convergence for two isolated single point calculations of (C-C equilibrium bond distance ). SQD configuration recovery calculations are performed in this case using sample size 5000 in 5 batches. Each column in Fig.7 shows the energy difference with HCI in log-scale (top row) and the number of configurations required (bottom row) in diagonalization subspace and search space of the respective methods against a common x-axis representing number of iterations. Column (A) and (B) shows calculations for C-C bond-length near equilibrium geometry and stretched geometry . PIGen-SQD convergence behaviour shows iteratively improved energy with more than one-order-of-magnitude better accuracy compared to SQD while keeping the diagonalization dimension and search space as low as possible. Quantitatively, even with a 75 reduction in the associated diagonalization subspace dimension for both the cases under consideration, energy accuracy of around 0.1 and 1 is achieved for equilibrium and stretched geometries compared to 5 and 10 for SQD, showing the superiority of PIGen-SQD under strong correlations.
Dissociation curve of with cc-pVDZ basis (with the frozen core approximation) is shown in Fig. 8. This is a challenging case with the associated symmetry space dimension . With sub-optimal shots, SQD struggles to identify the dominant configurations in this case as well and shows energy accuracy with SQD(5000) for all the geometries studied here compared to HCI. Contrarily PIGen-SQD provides consistently more accurate energy estimations () with only around 10 of the diagonalization cost of SQD.
@@ VI Conclusion and Future Outlook
In this paper, we propose PIGen-SQD, a novel QCSC workflow based on SQD framework that leverages a strategic orchestration of sampling efficiency of quantum computers, many-body intuitions and generative power of machine learning to provide better convergence with substantially reduced classical resource utilization. QCSC/ SQD utilizes the quantum computer as a sampling engine to generate configuration subspace in which the molecular Hamiltonian is projected and diagonalized. In the presence of quantum hardware noise, SQD requires a configuration recovery module, a heuristic technique that “corrects” the noise-affected samples with iterative filtering based upon the correct spin-particle number symmetry and average occupation probability. Accuracy of SQD relies upon the the degree of initial overlap with the target state provided by the sampled distribution and the capability of heuristic configuration recovery scheme to predict dominant configurations. However, the key challenge is that the state preparation on current quantum hardware does not necessarily guarantee high quality samples to provide this support to the target state and may require prohibitively huge number of shots which collectively adds a heuristic component to the SQD pipeline. To address this, here we introduce a physics-informed pre-processing which employs perturbative measures with efficiently designed implicit low rank tensor decompositions to identify a particular group of dominant configurations. These perturbative measure when combined with hardware samples, yields a substantially improved initial overlap to the target state providing a head-start for configuration recovery. This is followed by integration of generative ML models (RBM in this case) which upon training with this sample distribution, generates new configurations while exploring only the relevant sector of the Hilbert space of chemical relevance. Our numerical demonstrations with 27, 36, 45 and 52 qubit molecular simulations conducted on state-of-the-art IBM Heron R2 and R3 processors along with HPC resources suggest PIGen-SQD outperforms traditional SQD in both accuracy and efficiency, providing up to one-order of better energy accuracy with up to 90 reduction in the corresponding subspace dimension required for Hamiltonian diagonalization. Such accuracy with substantially reduced HPC resource requirements is observed even under strong electronic correlations where SQD struggles to predict chemically accurate results.
@@
From a near-future application perspective, PIGen-SQD can be integrated with other SQD variants such as Sample-based Krylov Quantum Diagonalization (SKQD)[47] or its more efficient variant using quantum stochastic drift protocol (SqDRIFT)[50] for guaranteed and faster convergence along with enhanced computational efficiency. Moreover, the current architecture have huge scope of technical improvements such as RBM can be replaced with more advanced neural network models such as transformers[78] which has recently been shown to provide unique advantages while dealing with complex molecular systems. With continued improvements in quantum-hardware capabilities and machine-learning components, PIGen-SQD presents a promising pathway toward scalable molecular simulations, advancing the field toward resolving several longstanding challenges in chemistry.
@@ Appendix A Algebraic Terms for Perturbative Calculations
The triply excited configurations arising from second-order perturbative corrections to the wave function are generated by the two terms in Eq. (22). Among these, the dominant computational bottleneck is the term , which involves contraction over a virtual orbital index. In practice, a given triply excited configuration can be generated through multiple contraction pathways. Consequently, the excitation manifolds produced by the and terms exhibit substantial overlap. Since the objective of the PIGen-SQD workflow is to provide a support to quantum-hardware-derived samples rather than to construct a complete perturbative expansion, the perturbative evaluation can be selectively restricted to those terms with more favorable computational scaling based on available classical HPC resources.
With the increase in the order of the perturbative correction, the number of terms contributing to a particular excitation manifold grows rapidly as shown in Eq. (23) for quadruples generations. In such cases also, the computations can be restricted to only a few terms if necessary as the PIGen-SQD workflow does not demand an exact calculation of all the terms in the expansion.